OpenStudy (anonymous):

A theater has 47 seats in first row, 50 in second row 53 in the third row. Can the number of seats in each row be modeled by arthmetric or geometric sequence. what type of sequence is present is ii area or geometric. a^n=, There are how many seats in row 19?

OpenStudy (anonymous):

50-47 = 3 extra seats in row 2, 53-50 = 3 extra seats in row 3... so we are adding 3 seats per row starting with row one witch has 47 seats....

OpenStudy (anonymous):

101seats

OpenStudy (anonymous):

hang on a sec, let me check

OpenStudy (anonymous):

thats what i got

OpenStudy (anonymous):

ok sow hat is a^n=

OpenStudy (anonymous):

so what does a^n=

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

Does a^n =101

OpenStudy (anonymous):

sorry, keeps freezing on me, i believe it would be area or arithmetic sequence and yes im thinking a^n in this case would be 101, but im not 100% certain on it

OpenStudy (anonymous):

thanks